3x^ -4x - 5 =0 use the quadratic formula

Answer by checkley71(8403)

3x^2-4x-5=0
x=(4+-sqrt[-4*2*-4*3-5])/2*3
x=(4+-sqrt[16+60])/6
x=(4+-sqrt76)/6
x=(4+-8.718)/6
x=(4+8.718)/6
x=12.718/6
x=2.1196 answer.
x=(4-8.718)/6
x=-4.718/6
x=-.786 answer.